OpenStudy (anonymous):

the sum of two intergers is at most -57. one integer is 23. what is the other integer? use x to represent the variable. follow this ex: x>25

OpenStudy (anonymous):

At most can also be stated as "less thank or equal to" for which the symbol is \[\le\]

OpenStudy (anonymous):

So, since 23 is a positive, do you need X to be positive, or negative?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

-57 is negative, so if 23 is positive ... X needs to be...

OpenStudy (anonymous):

negitive??

OpenStudy (anonymous):

Yes, but sadly, I have to go. :(

OpenStudy (anonymous):

If you wait a while I may be back, but i will tag someone who can help you

OpenStudy (anonymous):

k thx.

OpenStudy (anonymous):

@ParthKohli @jim_thompson5910 @Agent_Sniffles

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@dmezzullo

OpenStudy (anonymous):

As erdog said, "at most" can also be stated as "less than or equal to" for which the symbol is \[\le\]The missing addend is x. So the sum of x and 23, is at most -57. \[x + 23\le -57\]

OpenStudy (anonymous):

so x>-17?

OpenStudy (anonymous):

Not quite. You want to solve for x, the missing number. To do this simply subtract 23 from both sides. \[x + 23 -23\le -57-23\]\[x \le -57-23\]What is x?

OpenStudy (anonymous):

-80?

OpenStudy (anonymous):

Right! So x <= 80

OpenStudy (anonymous):

thx.

OpenStudy (anonymous):

no problem. (:

OpenStudy (anonymous):

can u help with another problem?

OpenStudy (anonymous):

I can ☺

OpenStudy (anonymous):

awesome her it is: miles worked a total of 38 hours last week. his earnings for the week were more than $228. what is his hourly rate of pay? follow this ex :x>25$

OpenStudy (anonymous):

Okay, so first, you want to divide 228 by 38

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Yup, but since it says "more than $228" what is the answer: x is __________than 6 ?

OpenStudy (anonymous):

less than

OpenStudy (anonymous):

more than instead

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

x>6

OpenStudy (anonymous):

x>$6 Good!

OpenStudy (anonymous):

If you think I was helpful, please press best response next to my answer ☺

OpenStudy (anonymous):

another q.?

OpenStudy (anonymous):

If you make/ post a new question, you could give another medal. ☺

OpenStudy (anonymous):

k